OpenStudy (anonymous):

A certain mountain has an elevation of 19,517ft. In 1916, the glacier on this peak covered 4 acres. By 2004 this glacier had melted to only 1 acre. a) Assume that this glacier melted at a constant rate each year. Find the yearly rate. b) Use the anser from part (a) to write a linear equation that gives the acreage A of this glacier t years past 1916.

OpenStudy (anonymous):

ok so in 2004-1916 years the glacier melted by 1 acre can you say what is the rate now ?

OpenStudy (anonymous):

Not sure what equation to use to find that rate.

OpenStudy (anonymous):

rate -> acres per year so in (2004-1916) = 88 years you lost 1 acre so the rate is (1/88) acres per year

OpenStudy (anonymous):

I get it now. Easy enough; just needed it broken down..

OpenStudy (anonymous):

ok .. can you answer (b) ?

OpenStudy (anonymous):

I can not figure it out.

OpenStudy (anonymous):

ok now you have the rate 1/88 but it is negative rate .. we are loosing glacier so we will put a minus sign in front of it. it will be our slope (call it m) and we have a "point" (t1,A1) = (1916,4) t for us is the year and A for us is the acres melted since they want the glacier after past 1916 we might call it 0 so our point is (0,4) so using in order to get the formula for the linear equation : A = A1 + m(t-t1) A(t) = 4 - (1/88) * (t-0) A(t) = 4 - t/88 here you see that after 88 years (at year 2004) you have A(88) = 3 which is like the data you have.

OpenStudy (anonymous):

I thought the slope would have to be used to find the equation to be used. Thanks it is so much easier when broken down into easier to understand terms :)